OpenStudy (anonymous):

0 is: rational irrational

OpenStudy (mertsj):

Remember that irrational numbers, in decimal form, go on and on forever and never repeat. Rational numbers can be expressed like this: \[\frac{3}{4}, \frac{23}{15}, \frac{0}{98} \]

OpenStudy (mertsj):

So where would you put 0?